OpenStudy (anonymous):

The area of a rectangular park is 3/5 square mile .the leght of the park is 7/8 mile what is the with of the park ?

OpenStudy (anonymous):

A = L W 3/5 = (7/8) (W) W = (3/5)(8/7)=24/35